Equal volumes,50.0 ml M ,of 0.54897solution of HCL and 0.9573 M NaOH solution having an initial temperature of 18.5 degree Celsius react in a coffee cup calorimeter.the resultant solution records a temperature of 22.9 degree Celsius.since the solution are mostly water,the solution are assumed to have a density of 1.0 g/mL and a specific heat of 4.184 J/g degree Celsius.The heat capacity of the calorimeter is 1.0 times 10 to power 1 J/degree Celsius.

Calculate the heat gained or released by the solution at constant pressure in J?


1
Expert's answer
2021-08-24T02:20:29-0400


ΔH=qp

0.9753 × 39.997 = 39.01g

Molar mass of HCl = 36.458

= 0.5489 × 50 × 36.458 = 1000.6g

Temperature change = 22.9 -18.5 = 4.1°C

grams of substance = 36.458 + 1000.6 = 1037.058 g

4.184 /1.037058 × 4.1

= 0.984J